The municipality wants to lay hexagonal paving stones around each flower bed. Using a flow chart and table, work out how many paving stones the municipality will need to pave around:2 flower beds, 16 flower beds and 32 flower beds. Write down the rule(also called a formula). Rule: ×4+2
1
Expert's answer
2020-06-30T18:44:39-0400

According to your rule 4x+2

we find out the value of paving stones by using this formula

For 2 flower beds , x=2

=4(2)+2

=10

For 16 flower beds ,x=16

=4(16)+2

=66

For 32 flower beds ,x=32

=4(32)+2

=130

TABLE